5. According to Boyles Law, if a system has an initial volume of 2.75 L which is at an pressure of 1.02 atm and the volume increases to 3.25 L. The final pressure of the gas should? increase B) decrease c) remain the same D) equal zero
Answer #1

Boyle's law can be mathematically expressed as PV= constant. P is pressure of gas and V is volume occupied by gas.

So, if initial Volume and Pressure are V1 and P1 and final volume and pressure are V2 and P2 respectively, then

P1*V1 = P2 * V2

So,

1.02 atm * 2.75L = P2 * 3.25L

So, P2 = 0.86atm

So, pressure decreases